# RacketVision: A Multiple Racket Sports Benchmark for Unified Ball and Racket Analysis

### v1

@@ Introduction
Racket sports, typically represented by badminton, tennis, and table tennis, have garnered widespread global participation and attract research for performance analysis (kulkarni_table_2023; ttrobo2024; gossard_tt3d_2025). These sports encompass structurally defined computer vision tasks, while presenting perception challenges due to the rapid motion of both the ball and players, as well as the complex human-object interactions inherent in racket-based gameplay. However, existing datasets (huang2019tracknet; sun2020tracknetv2; tarashima2023wasb) focus narrowly on ball tracking within a single sport at a time, falling short in two critical aspects: 1) They fail to leverage shared ball motion patterns across different sports. 2) Despite the racket being a central component, racket-specific annotations and analysis are lacking. This is crucial not only for sports analysi but also for complex neural avatar modeling (chen2024within; xu2025sequential; zhan2025r3; zhan2025towards). These shortcomings limit the development of comprehensive racket sports analysis methods.
To address this gap, we introduce RacketVision, a multiple racket sports benchmark for unified ball and racket analysis. RacketVision first expands the range of sports types for unified model training, aiming to uncover shared priors across racket sports. Specifically, it comprises a collection of 1,672 video clips (435,179 frames,  12,755 seconds) spanning badminton, tennis, and table tennis. In task design, RacketVision progressively proposes three tasks with corresponding annotations, enabling a more comprehensive decomposition of racket sport analysis. Beyond the existing ball tracking task, RacketVision defines racket keypoints and supports a novel racket pose estimation task. RacketVision further proposes an integrative task, ball trajectory prediction, empowering downstream applications, such as tactic analysis (wang2024tacticai; SGA-INTERACT), robotics (ttrobo2024; sci2025badmintonrobo), etc.
@@
Existing datasets in racket sports have primarily focused on ball tracking. As summarized in Table 1, while foundational datasets such as TrackNet (huang2019tracknet), TrackNetv2 (huang2019tracknet), and WASB (tarashima2023wasb) enabled deep learning approaches for single-sport ball tracking, they presented opportunities for expansion in terms of scene diversity, sport variety, and annotation scope. Building upon these efforts, RacketVision provides a significant leap in scale and diversity, featuring substantially more games and frames across three distinct sports. Critically, it introduces the first large-scale annotations for racket pose (R) in addition to ball positions (B), enabling novel multi-modal analysis beyond simple ball tracking.
@@
Research on sport analysis methods has evolved from the basic task of 2D ball tracking to more sophisticated tasks centered on humans and games (xia2024sportu; rao2024unisoccer; dong2024lucidaction; yang2024X; LearnableSMPLify). The development of robust 2D trackers has progressed from early CNN-based detectors (reno_convolutional_2018), to specialized architectures for small objects (jedrzejczak_deepball_2019), and data-efficient semi-supervised learning (vandeghen_semi-supervised_2022). The importance of this task is further underscored by large-scale benchmarks like SoccerNet (Deliege2020SoccerNetv2; cioppa_soccernet-tracking_2022). Recent studies have explored 3D trajectory and spin reconstruction from monocular videos (gossard_tt3d_2025; kienzle_towards_2025), hit anticipation (etaat_latte-mv_2025), and stroke recognition using trajectory data alone (kulkarni_table_2023). Racket-centric studies rely on specialized hardware like high-speed or stereo cameras (chen_visual_2013; gao_real-time_2019) or complex proxies like human keypoints to handle occlusions (zheng_method_2023). However, these advanced methods have been constrained by the lack of large-scale, unified benchmarks. RacketVision addresses this gap, providing a public benchmark to train and evaluate general-purpose models (jiang2023rtmpose; xu2022vitpose; carion2020end; cao2017realtime; yolo11_ultralytics) on these complex, interconnected analysis tasks, thereby lowering the barrier for future research.
@@
Settings. The task has two settings: (1) single-frame, using only the target frame $I_{t}$ to test static detection capabilities; and (2) multi-frame, using the target frame and five preceding frames $\{I_{t-5},\ldots,I_{t}\}$ to leverages temporal context for improved robustness against occlusions and motion blur (rozumnyi2021shape; zhong2022animation; zhong2023blur).
@@
  • RTMDet (lyu2022rtmdet): A state-of-the-art real-time object detection model, adapted for ball detection.
  • YOLO11 (yolo11_ultralytics): A state-of-the-art vision model for real-time object detection.
  • WASB (tarashima2023wasb): A strong baseline specifically designed for sports ball tracking, which internally incorporates background modeling.
  • TrackNetV3 (chen2023tracknetv3): A specialized heatmap-based network for tracking small, high-speed objects in sports, which can leverage temporal context from multiple frames.
@@
  • RTMPose (jiang2023rtmpose): A real-time top-down pose estimation model, optimized for keypoint detection in single-frame inputs. We adopt RTMDet (lyu2022rtmdet) as detector to generate bounding box.
@@
    • LSTM (hochreiter1997long): A 2-layer recurrent neural network that models temporal dependencies sequentially through a stateful, recurrent mechanism.
    • Transformer (vaswani2017attention): A 4-layer encoder-only model that captures global dependencies in the sequence in parallel via its self-attention mechanism.

### v3

@@ Introduction
Racket sports, typically represented by badminton, tennis, and table tennis, have garnered widespread global participation and attract research for performance analysis (Kulkarni et al. 2023; D’Ambrosio et al. 2024; Gossard et al. 2025). These sports encompass structurally defined computer vision tasks, while presenting perception challenges due to the rapid motion of both the ball and players, as well as the complex human-object interactions inherent in racket-based gameplay. However, existing datasets (Huang et al. 2019; Sun et al. 2020; Tarashima et al. 2023) focus narrowly on ball tracking within a single sport at a time, falling short in two critical aspects: 1) They fail to leverage shared ball motion patterns across different sports. 2) Despite the racket being a central component, racket-specific annotations and analysis are lacking. This is crucial not only for sports analysis but also for complex neural avatar modeling (Chen et al. 2024; Xu et al. 2025; Zhan et al. 2025a, b). These shortcomings limit the development of comprehensive racket sports analysis methods.
To address this gap, we introduce RacketVision, a multiple racket sports benchmark for unified ball and racket analysis. RacketVision first expands the range of sports types for unified model training, aiming to uncover shared priors across racket sports. Specifically, it comprises a collection of 1,672 video clips (435,179 frames,  12,755 seconds) spanning badminton, tennis, and table tennis. In task design, RacketVision progressively proposes three tasks with corresponding annotations, enabling a more comprehensive decomposition of racket sport analysis. Beyond the existing ball tracking task, RacketVision defines racket keypoints and supports a novel racket pose estimation task. RacketVision further proposes an integrative task, ball trajectory prediction, empowering downstream applications, such as tactic analysis (Wang et al. 2024; Yuchen et al. 2025b), robotics (D’Ambrosio et al. 2024; Ma et al. 2025), etc.
@@
Existing datasets in racket sports have primarily focused on ball tracking. As summarized in Table 1, while foundational datasets such as TrackNet (Huang et al. 2019), TrackNetv2 (Huang et al. 2019), and WASB (Tarashima et al. 2023) enabled deep learning approaches for single-sport ball tracking, they presented opportunities for expansion in terms of scene diversity, sport variety, and annotation scope. Building upon these efforts, RacketVision provides a significant leap in scale and diversity, featuring substantially more games and frames across three distinct sports. Critically, it introduces the first large-scale annotations for racket pose (R) in addition to ball positions (B), enabling novel multi-modal analysis beyond simple ball tracking.
@@
Research on sport analysis methods has evolved from the basic task of 2D ball tracking to more sophisticated tasks centered on humans and games (Xia et al. 2024; Rao et al. 2025; Dong et al. 2024; Yuchen et al. 2024, 2025a). The development of robust 2D trackers has progressed from early CNN-based detectors (Reno et al. 2018), to specialized architectures for small objects (Jedrzejczak et al. 2019), and data-efficient semi-supervised learning (Vandeghen et al. 2022). The importance of this task is further underscored by large-scale benchmarks like SoccerNet (Deliège et al. 2021; Cioppa et al. 2022). Recent studies have explored 3D trajectory and spin reconstruction from monocular videos (Gossard et al. 2025; Kienzle et al. 2025), hit anticipation (Etaat et al. 2025), and stroke recognition using trajectory data alone (Kulkarni et al. 2023). Racket-centric studies rely on specialized hardware like high-speed or stereo cameras (Chen et al. 2013; Gao 2019) or complex proxies like human keypoints to handle occlusions (Zheng et al. 2023). However, these advanced methods have been constrained by the lack of large-scale, unified benchmarks. RacketVision addresses this gap, providing a public benchmark to train and evaluate general-purpose models (Jiang et al. 2023; Xu et al. 2022; Carion et al. 2020; Cao et al. 2017; Jocher and Qiu 2024) on these complex, interconnected analysis tasks, thereby lowering the barrier for future research.
@@
Settings. The task has two settings: (1) single-frame, using only the target frame $I_{t}$ to test static detection capabilities; and (2) multi-frame, using the target frame and five preceding frames $\{I_{t-5},\ldots,I_{t}\}$ to leverages temporal context for improved robustness against occlusions and motion blur (Rozumnyi et al. 2021; Zhong et al. 2022, 2023).
@@
  • RTMDet (Lyu et al. 2022): A state-of-the-art real-time object detection model, adapted for ball detection.
  • YOLO11 (Jocher and Qiu 2024): A state-of-the-art vision model for real-time object detection.
  • WASB (Tarashima et al. 2023): A strong baseline specifically designed for sports ball tracking, which internally incorporates background modeling.
  • TrackNetV3 (Chen and Wang 2023): A specialized heatmap-based network for tracking small, high-speed objects in sports, which can leverage temporal context from multiple frames.
@@
  • RTMPose (Jiang et al. 2023): A real-time top-down pose estimation model, optimized for keypoint detection in single-frame inputs. We adopt RTMDet (Lyu et al. 2022) as detector to generate bounding box.
@@
    • LSTM (Hochreiter and Schmidhuber 1997): A 2-layer recurrent neural network that models temporal dependencies sequentially through a stateful, recurrent mechanism.
    • Transformer (Vaswani et al. 2017): A 4-layer encoder-only model that captures global dependencies in the sequence in parallel via its self-attention mechanism.